Solve
-3X +12+ 7X = 8+ 5X +4 -X

Answer by Cromlix(4381)

Hi there,
-3X +12+ 7X = 8+ 5X +4 -X
Collect all the X's on the left and
the numbers on the right.
Remember if you change the side with an X or
a number you change the sign
-3X + 7X - 5X + X = 8 + 4 - 12
0 = 0
Hope this helps :-)